# Solving Stochastic Orienteering Problems with Chance Constraints Using a GNN Powered Monte Carlo Tree Search

### v1

@@ Abstract
Leveraging the power of a graph neural network (GNN) with message passing, we present a Monte Carlo Tree Search (MCTS) method to solve stochastic orienteering problems with chance constraints. While adhering to an assigned travel budget the algorithm seeks to maximize collected reward while incurring stochastic travel costs. In this context, the acceptable probability of exceeding the assigned budget is expressed as a chance constraint. Our MCTS solution is an online and anytime algorithm alternating planning and execution that determines the next vertex to visit by continuously monitoring the remaining travel budget. The novelty of our work is that the rollout phase in the MCTS framework is implemented using a message passing GNN, predicting both the utility and failure probability of each available action. This allows to enormously expedite the search process. Our experimental evaluation shows that with the proposed method and architecture we manage to efficiently solve complex problem instances while incurring in moderate losses in terms of collected reward. Moreover, we demonstrate how the approach is capable of generalizing beyond the characteristics of the training dataset. The paper’s website, open-source code, and supplementary documentation can be found at ucmercedrobotics.github.io/gnn-sop.
@@ I Introduction
Recently, we proposed a novel solution to the stochastic orienteering problem using Monte Carlo Tree Search (MCTS), aimed at estimating both the collected reward and the probability of exceeding the allocated budget $B$ [2]. Previous work highlighted in [24] and confirmed through our own experiments demonstrates that the rollout phase of MCTS can consume up to 99% of the total planning time. Aiming at accelerating the search with the goal of deploying a planner in applications requiring real time response, in this work we investigate the potential of using a graph neural network (GNN) architecture to train a model capable of predicting rollout outcomes. A key innovation of our approach lies in simultaneously predicting two critical metrics simultaneously needed to solve SOP instances: the accumulated reward and the probability of budget constraint violations. To the best of our knowledge, this dual prediction has not been explored previously.
@@
  • we introduce a single network architecture capable of predicting values for a custom developed rollout procedure that outputs both action quality and failure probabilities;
@@
The rest of the paper is organized as follows. Selected related work is presented in Section II. SOP is formalized in Section III, and in the same section we briefly define the model that generates our training data. The core of our contribution is in Section IV where we describe the network we developed, as well as the feature attributes we identified to effectively train the networks. Extensive simulations detailing our findings are given in Section V, with conclusions and future work discussed in Section VI.
@@ II Related Literature
Early approaches to solve deterministic and stochastic orienteering problems relied on classical methods, often using heuristics due to the intrinsic computational complexity [8]. Notably, in [22] the authors propose an exact solution based on a mixed integer linear program formulation to solve the stochastic orienteering problem with chance constraints. This approach is however offline, i.e., it computes a solution before execution starts, and does not adapt on the fly. Our recent work [2] instead solves the problem using MCTS, designed to be an online method, whereby the vertices to visit are selected at runtime based on the remaining budget. This method, while competitive, devotes a significant amount of time to the rollout phase in MCTS search.
The practice of embedding policy and or value networks into MCTS has been notably used in [15] and in subsequent papers such as [28]. Even though these papers do not use GNNs, they provide evidence into being able to solve computationally expensive problems by reducing the simulation required to approximate a solution. They leverage the power of MCTS and approximate their rollout phases, among other improvements. Several recent approaches have explored the use of GNN-based solutions to solve NP-hard graph optimization problems. For example, [5, 24] focused on using GNNs to solve TSP and other graph optimization problems. While different from orienteering, solutions to TSP are of interest due to some similarities in the problem formulation. Both of these papers used a message passing framework with varying layer configuration. Importantly, [24] used MCTS to explore a limited horizon rather than the entire state space. In [11] the authors show that a similar implementation can be used to solve orienteering problems. In particular, a GNN-based approach is used in combination with a beam search method as done in [24]. However, none of these approaches address stochastic orienteering. Equally relevant, to the best of our knowledge, no methods combining GNNs and MCTS have been developed to solve the SOP with chance constraints, which is the focus of this contribution.
@@ III-A Stochastic Orienteering Problem with Chance Constraints
Stochastic Orienteering Problem with Chance Constraints (SOPCC) With the notation introduced above, let $10<P_{f}<$ 1 be an assigned failure bound. The SOPCC seeks to solve the following optimization problem:
@@
Remark: since the random variables associated with the edges are assumed to be known, generating random samples for the cost $C(P)$ of a path $P$ is straightforward and does not necessitate computing the probability density function of $C(P)$. This is consistent with the assumptions usually made when using MCTS to solve planning problems [9].
@@ III-B Monte Carlo Tree Search for SOPCC
In [2] we introduced a new approach based on MCTS to solve the SOPCC. Key to our new solution was the introduction of a novel criterion for tree search dubbed UCTF – Upper Confidence Bound for Tree with Failures. In this subsection we shortly review this solving method. We refer the reader to [2] for a deeper discussion of our method and to [4] for a general introduction to the MCTS methodology. Note that in [2] when deciding which vertex to add to the route, MCTS only considers the $K$-nearest neighbors to the current vertex. In the version implemented in this paper to train the GNN described in the next section, we remove this feature and instead consider all vertices. The reason is that in [2] we reduced the number of possibilities to expedite the online search, while here training is done offline. We can therefore consider a larger search space. As pointed out in [17] (chapter 8) a solution based on MCTS relies on the definition of four steps: selection, expansion, rollout and backup. In the following, we sketch how these can be customized for solving SOPCC.
@@ III-B1 Selection
Selection, also known as tree policy, guides the search from the root to a leaf node based on the value associated with the nodes in the tree. At every level, each child of the current node is evaluated using a metric, and the search then proceeds to the node with the highest value. Universal Confidence bound for Trees (UCT) is commonly used to attribute values to nodes [10] and defines the tree policy. UCT, however, aims at maximizing reward only and is therefore not suitable for use in problems with chance constraints where high reward could lead to violations of the constraint. To overcome this limit, in [2] we introduced UCTF, defined as follows
@@
where the term $(1-F[v_{j}])$ is new. In Eq. (1), $Q[v_{i}]$ is the estimate of how much reward will be collected by adding $v_{j}$ to the route, $F[v_{j}]$ is a failure probability estimate for a path going through $v_{j}$, $N[v_{j}]$ is the number of times $v_{j}$ has been explored and $N[v_{p}]$ is the number of times its parent has been explored. Critically, in our original approach both $Q$ and $F$ are estimated via rollout, while in this paper we will train a GNN to quickly predict these values. The term $(1-F[v_{j}])$ penalizes nodes with high estimates for the failure value. As in the classic MCTS method, the term $z$ balance exploration with exploitation.
@@ III-B2 Expansion
Expansion is the process of adding a node to the tree. In our implementation, this is implicitly obtained by assigning a UCTF value of infinity to vertices that are still unexplored (i.e., for which $N[v_{j}]$ is 0). This way, we ensure all children nodes of parent node are explored at least once before one of the siblings is selected again.
@@ III-B3 Rollout
Once a node is added, rollout (also called simulation) is used to estimate the amount of reward that will be collected by going through that node as well as its failure probability. This is done by running a baseline, handcrafted, heuristic policy multiple times. The $F$ value is estimated as the ratio of failed runs over total runs of the heuristic policy, while the $Q$ value is estimated as the average of the returned value limited to the successful runs. A critical aspect of this step that is the necessity of running the heuristic multiple times to numerically estimate $F$. In [2], the heuristic is a mix of random exploration and greedy search.
@@ III-B4 Backup
The backup step in [2] is unique because it propagates back not only the $Q$ value, but also the $F$ value. This is because for each action we must estimate not only the anticipated reward, $Q$, but also the probability, $F$, that the action will eventually result in a budget violation. To this end, two cases are considered by comparing the $F$ value of the parent with the $F$ value returned by the rollout. If the $F$ value for the parent node $v_{p}$ is $\leq P_{f}$, we check if the $F$ value of the newly expanded node is also below $P_{f}$. If it is, we then compare the reward estimate, $Q$. If the child node has a larger estimate with adding the reward of the parent node, we update the parent estimate. If not, we move on. In the case where the parent node has an $F>P_{f}$, we check if the child node $F$ estimate is less than the parent’s. If it is, we update the parent’s $F$ and $Q$ with $F$ being the child’s $F$ value and $Q$ being the child’s $Q$ plus the true, known reward of the parent node (see [2] for additional details).
@@ IV-A MCTS Value and Failure Network
The main novelty of the method we propose here is a single value/failure network architecture that predicts both the expected value, $Q$, and the failure probability, $F$, of an action, $v_{j}$, to be used in the tree policy UCTF formula from Eq. (1). Therefore, after appropriate training, the MCTS rollout phase with multiple simulations to estimate both $Q$ and $F$ is replaced by a single forward pass on each network, thus dramatically cutting the computation time. As a result, as explained later, when a node is considered for expansion all descendants are generated and evaluated at once. This is different from the classic MCTS approach where node descendants are added and evaluated one by one. This feature greatly extends the set of actions explored while not requiring more computational time.
@@ IV-B Message Passing Neural Network
Given that orienteering problem instances are defined over graphs, a message passing neural network (MPNN) [27, 26] is the natural choice to implement the value and failure networks. MPNNs are graph neural networks in which, at every iteration of learning, nodes of the graph share information with neighboring nodes to generate an $D$-dimensional embedding for each node or for the entire graph. In a typical message passing framework, the message update function encompasses node attributes as well as edge attributes, both of which are critical in the SOPCC problem space. The following equations show the general relationship between an embedding and a message within the framework [6], formulated in terms of the aggregation operation given by Eq. (2) (where the set $\mathcal{N}(v)$ is the set of nodes neighbors of $v$) and the update operation given by Eq. (3).
@@
Aggregation defines that a message $m_{v}^{t+1}$ for node $v$ at time $t+1$ is obtained by combining the embedding $h_{v}^{t}$ with all of its neighbors’ embeddings, $h_{w}^{t}$, at time $t$, as well as the attributes $e_{vw}$ of the edges connecting them. The update operation then creates the embedding $h_{v}^{t+1}$ for vertex $v$ at time $t+1$ by combining its previous embedding at time $t$ with the message at time $t+1$ obtained through aggregation. In both operations it is assumed that the data is transformed by a function, i.e., $M_{t}$ for aggregation and $U_{t}$ for update. In our implementation $M_{t}$ is defined as a matrix multiplication between edge attribute linear transformations and node embeddings. This was defined in [6] as
@@
The update function $U_{t}$ is defined as a Gated Recurrent Unit (GRU) layer introduced in [3]. Note that [6] states that the GRU will take as input $h_{v}^{t}$ and $h_{v}^{0}$ at every iteration, but we opt for the network from [13] that instead takes in $h_{v}^{t}$ and $h_{v}^{t-1}$. This design choice was made after experimentally observing that the latter gives better results than the former. Finally, we pass the output from the message passing layers into a multi-layer perceptron (MLP) to obtain the $Q$ and $F$ values. In our implementation, we use three message passing layers, i.e., the maximum value for $t$ is three. Critical to the approach are the node attributes to be used for $h$ at time 00, as well as the edge attributes $e_{vw}$ prior to linearization. These attributes will be defined in the remainder of this section. In our use case, we create two different networks using this architecture, one for $Q$ and one for $F$ [^1] The complete GNN architecture as well as relevant training documentation can be found at ucmercedrobotics.github.io/gnn-sop.. We have two separate networks since the MCTS training data is best captured by two different activation functions for $Q$ and $F$, We select a linear activation function for the $Q$ value and a sigmoid for $F$. The reasoning is that $Q$ predicts a positive, but potentially unbounded scalar value while $F$ predicts a probability between 0 and 1. We also hypothesized that the networks would train differently on both types of labels, so we opted to separate them. Indeed, preliminary experiments showed that our loss was reduced when separating the models completely instead of having one model with two output activation functions or two output MLPs. We selected mean squared error (MSE) as the loss function for both $Q$ and $F$ outputs, as the task at hand is regression.
@@ IV-C Attributes
For each node in the graph, we introduce an eight dimensional vector of attributes that will be used to define $h_{v}$ Eq. (2) for $t=0$. Each node has the following attributes:
@@
Our choice for these attributes starts from [24] where a similar approach was used to solve the traveling salesman problem (TSP). Accordingly, the first three attributes give the model the spatial awareness that proved to work well in TSP. However, since SOPCC has more requirements that TSP, we added additional attributes such as reward and residual budget owing to the fact that the nodes are not all uniformly important and in general the agent will not be able to visit each one. Intuition says that more spatial and temporal information will aid the model in learning the greedy heuristic we described in Section III-B. Indeed, as we will show in Section V-B through an ablation study, these additions were critical in performing comparably to previous methods. As for edge attributes $e_{vw}$, we set the single attribute to the Euclidean distance between nodes $v$ and $w$. Note that the edge attributes do not include any information about the length variability $d(e_{v,w})$. This choice was made after preliminary tests showed no conclusive performance improvement and a reduction in model parameters, allowing for a smaller training dataset and, ultimately, reducing training time.
In selecting our model, we started from the literature discussed in Section II as well as a test case comparison. We trained both a MPNN [6] and a graph attention network (GAT) [23]. Our starting hypothesis was that an attention network might be able to identify more valuable nodes through the attention mechanism. However, it turned out that the performance was a bit worse when training the GAT. To be specific, the average reward computed was roughly 10% worse in our largest model, with only  50ms gains in computational time through the entire MCTS pipeline, inference included. This trade off was, therefore, not substantial enough for selecting the GAT. Consequently, we selected the neural message passing model from [6, 13] for its performance and stability. Notably, the pre-built model from [13] includes a skip level connection from the initial node attributes to post-message passing layers. We removed this skip layer to be consistent with [6] and also because we observed it did not yield better results.

### v2

@@ Abstract
Leveraging the power of a graph neural network (GNN) with message passing, we present a Monte Carlo Tree Search (MCTS) method to solve stochastic orienteering problems with chance constraints. While adhering to an assigned travel budget, the algorithm seeks to maximize collected reward while incurring stochastic travel costs. In this context, the acceptable probability of exceeding the assigned budget is expressed as a chance constraint. Our MCTS solution is an online and anytime algorithm, alternating planning and execution, that determines the next vertex to visit by continuously monitoring the remaining travel budget. The novelty of our work is that the rollout phase in the MCTS framework is implemented using a message-passing GNN, predicting both the utility and failure probability of each available action. This allows to enormously expedite the search process. Our experimental evaluation shows that with the proposed method and architecture, we manage to efficiently solve complex problem instances while incurring moderate losses in terms of collected reward. Moreover, we demonstrate how the approach is capable of generalizing beyond the characteristics of the training dataset. The paper’s website, open-source code, and supplementary documentation can be found at ucmercedrobotics.github.io/gnn-sop.
@@ I Introduction
Recently, we proposed a novel solution to the stochastic orienteering problem using Monte Carlo Tree Search (MCTS), aimed at estimating both the collected reward and the probability of exceeding the allocated budget $B$ [2]. Previous work highlighted in [24] and confirmed through our own experiments demonstrates that the rollout phase of MCTS can consume up to 99% of the total planning time. Aiming to accelerate the search with the goal of deploying a planner in applications requiring real-time response, in this work we investigate the potential of using a graph neural network (GNN) architecture to train a model capable of predicting rollout outcomes. A key innovation of our approach lies in simultaneously predicting two critical metrics needed to solve SOP instances: the accumulated reward and the probability of budget constraint violations. To the best of our knowledge, this dual prediction has not been explored previously.
@@
  • we introduce a single architecture capable of predicting values for a custom developed rollout procedure that can output action quality and failure probability;
@@
The rest of the paper is organized as follows. Selected related work is presented in Section II. SOP is formalized in Section III, and in the same section we briefly define the model that generates our training data. The core of our contribution is in Section IV, where we describe the network we developed, as well as the feature attributes we identified to effectively train the networks. Extensive simulations detailing our findings are given in Section V, with conclusions and future work discussed in Section VI.
@@ II Related Literature
Early approaches to solve deterministic and stochastic orienteering problems relied on classical methods, often using heuristics due to the intrinsic computational complexity [8]. Notably, in [22] the authors propose an exact solution based on a mixed-integer linear program formulation to solve the stochastic orienteering problem with chance constraints. This approach is however offline, i.e., it computes a solution before execution starts and does not adapt on the fly. Our recent work [2] instead solves the problem using MCTS, designed to be an online method, whereby the vertices to visit are selected at runtime based on the remaining budget. This method, while competitive, devotes a significant amount of time to the rollout phase in MCTS search.
The practice of embedding policy and/or value networks into MCTS has been notably used in [15] and in subsequent papers such as [28]. Even though these papers do not use GNNs, they provide evidence of being able to solve computationally expensive problems by reducing the simulation required to approximate a solution. They leverage the power of MCTS and approximate their rollout phases, among other improvements. Several recent approaches have explored the use of GNN-based solutions to solve NP-hard graph optimization problems. For example, [5, 24] focused on using GNNs to solve TSP and other graph optimization problems. While different from orienteering, solutions to TSP are of interest due to some similarities in the problem formulation. Both of these papers used a message-passing framework with varying layer configurations. Importantly, [24] used MCTS to explore a limited horizon rather than the entire state space. In [11], the authors show that a similar implementation can be used to solve orienteering problems. In particular, a GNN-based approach is used in combination with a beam search method, as done in [24]. However, none of these approaches address stochastic orienteering. Equally relevant, to the best of our knowledge, no methods combining GNNs and MCTS have been developed to solve the SOP with chance constraints, which is the focus of this contribution.
@@ III-A Stochastic Orienteering Problem with Chance Constraints
Stochastic Orienteering Problem with Chance Constraints (SOPCC) With the notation introduced above, let $0<P_{f}<1$ be an assigned failure bound. The SOPCC seeks to solve the following optimization problem:
@@
Remark: Since the probability density functions of the random variables associated with the edges are assumed to be known, generating random samples for the cost $C(P)$ of a path $P$ is straightforward and does not necessitate computing the probability density function of $C(P)$. This is consistent with the assumptions usually made when using MCTS to solve planning problems [9].
@@ III-B Monte Carlo Tree Search for SOPCC
In [2], we introduced a new approach based on MCTS to solve the SOPCC. Key to our new solution was the introduction of a novel criterion for tree search dubbed UCTF – Upper Confidence Bound for Tree with Failures. In this subsection, we briefly review this solving method. We refer the reader to [2] for a deeper discussion of our method and to [4] for a general introduction to the MCTS methodology. Note that in [2], when deciding which vertex to add to the route, MCTS only considers the $K$-nearest neighbors to the current vertex. In the version implemented in this paper to train the GNN described in the next section, we remove this feature and instead consider all vertices. The reason is that in [2] we reduced the number of possibilities to expedite the online search, while here training is done offline. We can therefore consider a larger search space. As pointed out in [17] (chapter 8), a solution based on MCTS relies on the definition of four steps: selection, expansion, rollout, and backup. In the following, we sketch how these can be customized for solving SOPCC.
@@ III-B1 Selection
Selection, also known as the tree policy, guides the search from the root to a leaf node based on the value associated with the nodes in the tree. At every level, each child of the current node is evaluated using a metric, and the search then proceeds to the node with the highest value. Universal Confidence Bound for Trees (UCT) is commonly used to attribute values to nodes [10] and defines the tree policy. UCT, however, aims at maximizing reward only and is therefore not suitable for use in problems with chance constraints where high reward could lead to violations of the constraint. To overcome this limit, in [2] we introduced UCTF, defined as follows
@@
where the term $(1-F[v_{j}])$ is new. In Eq. (1), $Q[v_{i}]$ is the estimate of how much reward will be collected by adding $v_{j}$ to the route, $F[v_{j}]$ is a failure probability estimate for a path going through $v_{j}$, $N[v_{j}]$ is the number of times $v_{j}$ has been explored, and $N[v_{p}]$ is the number of times its parent has been explored. Critically, in our original approach both $Q$ and $F$ are estimated via rollout, while in this paper we will train a GNN to quickly predict these values. The term $(1-F[v_{j}])$ penalizes nodes with high estimates for the failure value. As in the classic MCTS method, the term $z$ balances exploration with exploitation.
@@ III-B2 Expansion
Expansion is the process of adding a node to the tree. In our implementation, this is implicitly obtained by assigning a UCTF value of infinity to vertices that are still unexplored (i.e., for which $N[v_{j}]$ is 0). This way, we ensure all children nodes of a parent node are explored at least once before one of the siblings is selected again.
@@ III-B3 Rollout
Once a node is added, rollout (also called simulation) is used to estimate the amount of reward that will be collected by going through that node as well as its failure probability. This is done by running a baseline, handcrafted, heuristic policy multiple times. The $F$ value is estimated as the ratio of failed runs over total runs of the heuristic policy, while the $Q$ value is estimated as the average of the returned value limited to the successful runs. A critical aspect of this step is the necessity of running the heuristic multiple times to numerically estimate $F$. In [2], the heuristic is a mix of random exploration and greedy search.
@@ III-B4 Backup
The backup step in [2] is unique because it propagates back not only the $Q$ value, but also the $F$ value. This is because for each action we must estimate not only the anticipated reward, $Q$, but also the probability, $F$, that the action will eventually result in a budget violation. To this end, two cases are considered by comparing the $F$ value of the parent with the $F$ value returned by the rollout. If the $F$ value for the parent node $v_{p}$ is $\leq P_{f}$, we check if the $F$ value of the newly expanded node is also below $P_{f}$. If it is, we then compare the reward estimate, $Q$. If the child node has a larger estimate when adding the reward of the parent node, we update the parent estimate. If not, we move on. In the case where the parent node has an $F>P_{f}$, we check if the child node $F$ estimate is less than the parent’s. If it is, we update the parent’s $F$ and $Q$ with $F$ being the child’s $F$ value and $Q$ being the child’s $Q$ plus the true, known reward of the parent node (see [2] for additional details).
@@ IV-A MCTS Value and Failure Network
The main novelty of the method we propose here is a single value/failure network architecture that can be trained to predict either expected value, $Q$, or failure probability, $F$, of an action, $v_{j}$, to be used in the tree policy UCTF formula from Eq. (1). Therefore, after appropriate training, the MCTS rollout phase with multiple simulations to estimate both $Q$ and $F$ is replaced by a single forward pass on each network, thus dramatically cutting the computation time. As a result, as explained later, when a node is considered for expansion, all descendants are generated and evaluated at once. This is different from the classic MCTS approach where node descendants are added and evaluated one by one. This feature greatly extends the set of actions explored while not requiring more computational time.
@@ IV-B Message Passing Neural Network
Given that orienteering problem instances are defined over graphs, a message passing neural network (MPNN) [27, 26] is the natural choice to implement the value and failure networks. MPNNs are graph neural networks in which, at every iteration of learning, nodes of the graph share information with neighboring nodes to generate a $D$-dimensional embedding for each node or for the entire graph. In a typical message passing framework, the message update function encompasses node attributes as well as edge attributes, both of which are critical in the SOPCC problem space. The following equations show the general relationship between an embedding and a message within the framework [6], formulated in terms of the aggregation operation given by Eq. (2) (where the set $\mathcal{N}(v)$ is the set of nodes neighboring $v$) and the update operation given by Eq. (3).
@@
Aggregation defines that a message $m_{v}^{t+1}$ for node $v$ at time $t+1$ is obtained by combining the embedding $h_{v}^{t}$ with all of its neighbors’ embeddings, $h_{w}^{t}$, at time $t$, as well as the attributes $e_{vw}$ of the edges connecting them. The update operation then creates the embedding $h_{v}^{t+1}$ for vertex $v$ at time $t+1$ by combining its previous embedding at time $t$ with the message at time $t+1$ obtained through aggregation. In both operations, it is assumed that the data is transformed by a function, i.e., $M_{t}$ for aggregation and $U_{t}$ for update. In our implementation, $M_{t}$ is defined as a matrix multiplication between edge attribute linear transformations and node embeddings. This was defined in [6] as
@@
The update function $U_{t}$ is defined as a Gated Recurrent Unit (GRU) layer introduced in [3]. Note that [6] states that the GRU will take as input $h_{v}^{t}$ and $h_{v}^{0}$ at every iteration, but we opt for the network from [13] which instead takes in $h_{v}^{t}$ and $h_{v}^{t-1}$. This design choice was made after experimentally observing that the latter gives better results than the former. Finally, we pass the output from the message passing layers into a multi-layer perceptron (MLP) to obtain the $Q$ and $F$ values. In our implementation, we use three message passing layers, i.e., the maximum value for $t$ is three. Critical to the approach are the node attributes to be used for $h$ at time 0, as well as the edge attributes $e_{vw}$ prior to linearization. These attributes will be defined in the remainder of this section. In our use case, we create two different networks using this architecture, one for $Q$ and one for FF [^1] The complete GNN architecture as well as relevant training documentation can be found at ucmercedrobotics.github.io/gnn-sop.. We have two separate networks since the MCTS training data is best captured by two different activation functions for $Q$ and $F$, We select a linear activation function for the $Q$ value and a sigmoid for $F$. The reasoning is that $Q$ predicts a positive, but potentially unbounded scalar value while $F$ predicts a probability between 0 and 1. We also hypothesized that the networks would train differently on both types of labels, so we opted to separate them. Indeed, preliminary experiments showed that our loss was reduced when separating the models completely instead of having one model with two output activation functions or two output MLPs. We selected mean squared error (MSE) as the loss function for both $Q$ and $F$ outputs, as the task at hand is regression.
@@ IV-C Attributes
For each node in the graph, we introduce an eight dimensional vector of attributes that will be used to define $h_{v}$ in Eq. (2) for $t=0$. Each node has the following attributes:
@@
Our choice for these attributes is inspired by [24], where a similar approach was used to solve the TSP. Accordingly, the first three attributes provide the model with spatial awareness, which proved effective in TSP setting. However, since SOPCC has more constraints than TSP, we added additional attributes such as reward and residual budget. This is due to the fact that the nodes are not all uniformly important, and in general the agent will not be able to visit every node. Intuitively, incorporating more spatial and temporal information helps the model learn the greedy heuristic described in Section III-B. Indeed, as we will show in Section V-B through an ablation study, these additions were critical to achieving performance comparable to previous methods. For edge attributes $e_{vw}$, we set the single attribute to the Euclidean distance between nodes $v$ and $w$. Note that the edge attributes do not include any information about the length variability $d(e_{v,w})$. This choice was based on preliminary tests, which showed no conclusive performance improvement from including this information. Additionally, omitting it reduces the number of model parameters, enabling a smaller training dataset and reducing training time.
In selecting our model architecture, we drew from the literature reviewed in Section II as well as a direct comparison using test cases. We trained both a Message Passing Neural Network (MPNN) [6] and a Graph Attention Network (GAT) [23]. Our initial hypothesis was that the attention mechanism in the GAT might help identify more valuable nodes by assigning higher weights to more promising neighbors. However, in practice, the GAT underperformed slightly compared to the MPNN. Specifically, the average reward achieved by the GAT was approximately 10% lower than that of our largest MPNN model, with only a marginal computational speedup of about 50 ms over the entire MCTS pipeline, including inference. This trade-off was not significant enough to justify selecting the GAT. Therefore, we opted for the neural message passing model presented in [6, 13], prioritizing its performance and training stability. Notably, the implementation from [13] includes a skip connection from the initial node attributes to the post-message-passing layers. We removed this skip layer to align more closely with the original design in [6] and because it did not result in any performance improvements in our experiments.